(1+yz)dx + z(z-x)dy - ( 1 + xy ) dz = 0 verify that the pfaffian differential equation are intergrable and find corresponding integral


1
Expert's answer
2021-07-27T07:50:47-0400

This is a Pfaffian differential equation in three variables and we must verify  its integrabilty and determine its primitive.


"(1+yz)dx+z(z\u2212x)dy\u2212(1+xy)dz=0"

The necessary and sufficient condition for integrability is


"X\u22c5curlX=0"

"X=(1+yz,z^2-xz,\u22121\u2212xy),"

so that


"\nabla \times X=\begin{vmatrix}
 i & j & k \\
 \dfrac{\partial}{\partial x} & \dfrac{\partial}{\partial y} & \dfrac{\partial}{\partial z}\\
 1+yz & z^2-xz & -1-xy \\
\end{vmatrix}"


"=i(-x-2z+x)-j(-y-y)+k(-z-z)"

"=-2zi+2yj-2zk"

"=(1+yz,z^2-xz,−1−xy)\cdot(-2z, 2y, -2z)"

"=-2z-2yz^2+2yz^2-2xyz+2z+2xyz=0"

Thus, the given equation is integrable.


Solving by Inspection


"(1+yz)dx+z(z\u2212x)dy\u2212(1+xy)dz=0"


"(1+yz)dx-(1+yz)dz+(1+yz)dz"

"\u2212(1+xy)dz+z(z\u2212x)dy=0"

"-(1+yz)d(z-x)+y(z-x)dz+z(z-x)dy=0"

"-(1+yz)d(z-x)+(z-x)d(1+yz)=0"

"\dfrac{d(1+yz)}{1+yz}-\dfrac{d(z-x)}{z-x}=0"

Integrating both sides, we have


"\ln(1+yz)-\ln(z-x)=\ln C"

"1+yz=C(z-x)"

"z=\dfrac{1+Cx}{C-y}"

is a solution to the Pfaffian differential equation.